# Quantum accessible information and classical entropy inequalities

### v2

@@ 5.2 Proof of theorem 3
It is easy to see that for $\alpha$ and $p$ such that $10<\alpha\leq 1/2\leq 1-\alpha\leq p<$ 1 one has $\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\geq 0$ and ${\sqrt{(1-p)(1-\alpha)}-\sqrt{p\alpha}\leq 0}$, therefore, $\tilde{\mu}_{0}(p,\alpha)\leq 0$.
The initial inequality (66) for $k_{1}c_{1}^{2}=t$ is equivalent to
$-t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}\pm\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m),$ (90)
if ±$\pm c_{2}\geq 0$.
Since $\tilde{\mu}_{0}(p,\alpha)\leq 0$,
$\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}+\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m)\leq$
$\leq\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m),$
@@
Proof of lemma 5.
So, for $\frac{m-1}{m}\leq p<1$ we have
$-\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(1-1/m)=-\dfrac{m-2}{m}\log(m-1)+\log m.$
For $p(m)\leq p<1$ we have
$-\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(p(m))=1,$
because the condition (71) is equivalent to $t>1$, under which $\tilde{\mu}_{1}\equiv-c_{1}=-1$.
@@
$-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}\geq\tilde{\mu}_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}-\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,\alpha).$
@@
Collecting similar terms,
$f(t):=2\sqrt{\beta(1-\beta)}\Bigl{(}\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)\Bigl{)}\sqrt{t(1-t)}-(1-2\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)}t+$
@@
The function $f(t)$ has the form $A\sqrt{t(1-t)}+Bt+C$ with $A\leq 0$ and $B\leq 0$, then the unique minimum point $t_{*}\in[1/2,1]$. it has the value $t_{*}=\dfrac{1}{2}\Bigl{(}1-\dfrac{B}{\sqrt{A^{2}+B^{2}}}\Bigl{)}$ and the minimum of the function $f(t)$ is equal to
$f(t_{*})=\dfrac{B}{2}+C-\dfrac{1}{2}\sqrt{A^{2}+B^{2}},$
@@
×$\times\Bigl{(}(2p-1)-2(p+\alpha-1)\Bigl{)}=$
@@
×$\times\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}$
@@
Let us denote $u(t)=4t(1-t)$. The inequality (91) can be rewritten in the form
@@
$t\mapsto g(t)=\dfrac{u(\beta)t}{u(\alpha)}+\dfrac{(1-u(\beta))(1-t)}{1-u(\alpha)}$
has non-negative derivative $\dfrac{u(\beta)}{u(\alpha)}-\dfrac{1-u(\beta)}{1-u(\alpha)}=\dfrac{u(\beta)-u(\alpha)}{u(\alpha)(1-u(\alpha))}$ and $g(u(\alpha))=1$. Therefore, the inequality (91) holds, and lemma has been proven. ∎

### v3

@@ 5.2 The case of theorem 3
It is easy to see that for $\alpha$ and $p$ such that $0<\alpha\leq 1/2\leq 1-\alpha\leq p<1$ one has $\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\geq 0$ and $\sqrt{(1-p)(1-\alpha)}-\sqrt{p\alpha}\leq 0$}, therefore, $\tilde{\mu}_{0}(p,\alpha)\leq 0$.
The initial inequality (67) for $k_{1}s_{1}^{2}=t$ is equivalent to
$t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m),$ (92)
@@
$\tilde{f}(t)\equiv-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigg{(}\tilde{\mu}_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}-\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,\alpha)\Bigg{)}\geq 0.$
@@
Collecting similar terms and using notation of $f(t)$ from subsection 5.1,
$\tilde{f}(t)-f(t)=2\sqrt{\beta(1-\beta)}\Bigl{(}\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)\Bigl{)}\sqrt{t(1-t)}-(1-2\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)}t+$
@@
The function $(\tilde{f}-f)(t)$ has the form $A\sqrt{t(1-t)}+Bt+C$ with $A\leq 0$ and $B\leq 0$, then the unique minimum point $t_{*}\in[1/2,1]$. It has the value $t_{*}=\dfrac{1}{2}\Bigl{(}1-\dfrac{B}{\sqrt{A^{2}+B^{2}}}\Bigl{)}$ and the minimum of the function $(\tilde{f}-f)(t)$ is equal to
$(\tilde{f}-f)(t_{*})=\dfrac{B}{2}+C-\dfrac{1}{2}\sqrt{A^{2}+B^{2}},$
@@
$\times\Bigl{(}(2p-1)-2(p+\alpha-1)\Bigl{)}=$
@@
$\times\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}$
@@
Let us denote u(t)=4t(1t)u(t)=4t(1-t). The inequality (93) can be rewritten in the form
@@
$t\mapsto W(t)=\dfrac{u(\beta)t}{u(\alpha)}+\dfrac{(1-u(\beta))(1-t)}{1-u(\alpha)}$
has non-negative derivative $\dfrac{u(\beta)}{u(\alpha)}-\dfrac{1-u(\beta)}{1-u(\alpha)}=\dfrac{u(\beta)-u(\alpha)}{u(\alpha)(1-u(\alpha))}$ and $W(u(\alpha))=1$}. Therefore, the inequality (93) holds, and lemma has been proven. ∎